On an RC circuit where C=1/400F , R=600Ω , and E(t) = 4 if 0<= t <3, -4 if 3<= t <= 6, 0 if 6<=t
the units of E (t) are Volts. Find the charge in Colulombs on the Capacitor at time t = 9 seconds. Take q (0) = 0.
Answer #1

XRC) at t=g 9.97X163 C a=2. (1-e (20 = CVo = 4 x 4 = 0.01C) q=0.0 (1-e - 8x400) q=0.01 (1-é AS